40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Aug 31 13:33:02 2010, tara, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
    Reply  Tue Aug 31 13:40:00 2010, Frank, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
       Reply  Tue Aug 31 14:50:53 2010, tara, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
          Reply  Tue Aug 31 14:52:41 2010, Frank, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
          Reply  Tue Aug 31 15:04:51 2010, Frank, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
             Reply  Tue Aug 31 16:15:51 2010, tara, Notes, Computers, dead channel, C3:PSL-ACAV_TEMPAVG 
Message ID: 330     Entry time: Tue Aug 31 16:15:51 2010     In reply to: 324
Author: tara 
Type: Notes 
Category: Computers 
Subject: dead channel, C3:PSL-ACAV_TEMPAVG 

Quote:

be carefull with the data you are taking right now. it's wrong for your projection as the power fluctuations are different when locking only the ACAV using the AOM. The largest contributor might be the pointing from the AOM itself, which is different if the laser isn't locked to the other cavityat the same time.

Why don't you use the new fast channels you have hooked up last week? And don't forget to change the names of those :-)

Quote:

Quote:

looks like some fault of the database. /usr1/epics/psl/db/acav.db does not contain the correct entry. check rcav.db and copy the record for "C3:PSL-RCAV_TEMPAVG" into acav.db. Then simply change "RCAV" into "ACAV" everywhere for this record. Also change the setpoint for the ACAV in the startup.cmd file to 37.3 and the ACAV-heater value to 4.653. Those are the latest values when both where locked for several hours. Reset the crate again.

Quote:

Yesterday, I reset the PSL crate behind the SUN computer, but the channel C3:PSL-ACAV_TEMPAVG is stil inactive.

 

 

 Will do in a moment, I'm taking data from ACAV for now just to compare with yesterday results.

 I didn't disabled the loop when I measured it, and yeah it looks bad. I think I'll just try to lock both cavities for now.

I'll connect the signals from PDs behind Rcav and Acav to the new fast channels(32k).

About changing the channels' names, I asked DMASS to help and he suggested to use them as they are for now because of the risk of screwing the system up from a typo.

ELOG V3.1.3-